OpenStudy (theedwardsfamily):

MEDAL WILL BE REWARDED! Solve the equation. x + 1/6 = 6 A. x = –6 1/6 B. x = –5 5/6 C. x = 5 5/6 D. x = 6 1/6

OpenStudy (anonymous):

5 5/6 + 1/6 = 6 or 35/6 + 1/6 = 6 x = 5 5/6 or x = 35/ 6 and if you need decimal x = 5.83333333333333333333333333333333333333... round up 5.83

OpenStudy (anonymous):

I think it will because 5 and 5/6 plus 1/6 equals 6

OpenStudy (anonymous):

5 5/6

OpenStudy (anonymous):

i bet i am not helping or am i

OpenStudy (anonymous):

x=35/6

OpenStudy (anonymous):

well you do 6 - 1/6 which is 5 5/6

OpenStudy (anonymous):

so C

OpenStudy (anonymous):

because by putting 35/6 instead of x and adding it to 1/6 we get the answer 6

OpenStudy (anonymous):

Sari as long as you try it counts :)

OpenStudy (theedwardsfamily):

thanks! but who should I medal? :/